# Implementing feature binding through dendritic networks of a single neuron

### v1

@@ Abstract
A single neuron receives an extensive array of synaptic inputs through its dendrites, raising the fundamental question of how these inputs undergo integration and summation, culminating in the initiation of spikes in the soma. Experimental and computational investigations have revealed various modes of integration operations that include linear, superlinear, and sublinear summation. Interestingly, distinct neuron types exhibit diverse patterns of dendritic integration contingent upon the spatial distribution of dendrites. The functional implications of these specific integration modalities remain largely unexplored. In this study, we employ the Purkinje cell as a model system to investigate these intricate questions. Our findings reveal that Purkinje cells (PCs) generally exhibit sublinear summation across their expansive dendrites. The degree of sublinearity is dynamically modulated by both spatial and temporal input. Strong sublinearity necessitates that the synaptic distribution in PCs be globally scattered sensitive, whereas weak sublinearity facilitates the generation of complex firing patterns in PCs. Leveraging dendritic branches characterized by strong sublinearity as computational units, we demonstrate that a neuron can adeptly address the feature-binding problem. Collectively, these results offer a systematic perspective on the functional role of dendritic sublinearity, providing inspiration for a broader understanding of dendritic integration across various neuronal types.
@@ I Introduction
Neurons, the foundational computational entities within the intricate brain system, intricately process signal inputs received through their dendritic spine. Conventional understanding of dendritic function posits that these structures receive synaptic input and subsequently convey it to the soma, thus initiating the generation of action potentials or spikes. This proposition suggests that a single neuron possesses computational capabilities to execute information processing through an extensive array of dendritic branches. However, the underlying integrating principles governing the diverse array of dendrites and neuron types remain elusive. In recent decades, experimental studies have facilitated laborious explorations of the integration of synaptic inputs along the dendritic domains of individual neurons, shedding light on the physiological basis of dendrite function [1, 2, 3, 4]. Dendrites have been shown to function as autonomous information processing units, capable of performing local computations, thus conferring substantial computational capabilities on single neurons [5, 6, 7, 8, 9, 10]. Furthermore, these insights have significantly contributed to our understanding of how neurons engage in communication with one another and the resultant impact on behavior [11, 12, 13, 14].
@@
The principal function of dendrites lies in the intricate integration and processing of synaptic inputs [6, 15]. Investigations focusing on pyramidal cells reveal that dendritic integration is characterized by superlinear dynamics, attributed to the activation of voltage-dependent channels and NMDARs [12, 16, 17, 18]. In particular, this superlinear integration has been observed primarily in cortical neurons [11, 19, 20], while sublinear summation has been documented in cerebellar interneurons  [21, 22]. The manifestation of dendritic nonlinearity allows neurons to function as feature detectors [14, 13], exerting influence on various aspects of brain function, including whisker sensation [23], orientation selectivity [24], sensory perception [25, 26], sensory-motor integration [23], and memory encoding [27, 28]. Furthermore, nonlinear dendritic integration is postulated to augment the computational capacity of neurons [29, 30], enabling them to compute linearly non-separable functions [31, 32]. The incorporation of nonlinear dendritic mechanisms into artificial neural networks is proposed to bring neurons in comparative alignment with their biological counterparts [33, 34, 28, 5]. The benefit of installing dendritic nonlinearity into artificial neural networks not only reduces power consumption and enhances accuracy [35, 36] but also mitigates communication costs within neural networks [37].
@@
The majority of investigations into dendritic nonlinearity have predominantly centered on cortical neurons. However, there is a substantial gap in our understanding of the dendritic nonlinearity inherent in the most complex neuronal entities, namely cerebellar Purkinje cells (PCs) [38]. The intricate dendritic structures of PCs contribute to amplification of the computational repertoire within single neurons. PCs, characterized by receiving thousands of synaptic inputs in a location-dependent manner, exhibit a distinctive synaptic organization, with the distal region of the dendritic tree receiving excitatory input from parallel fibers originating from granule cells [39]. On the contrary, the proximal portion of the dendritic tree is innervated by the input of climbing fibers [40]. The interplay between synaptic input and dendritic nonlinear properties produces highly intricate input-output functions [34]. Therefore, it is imperative to understand whether passive cable properties contribute to location-dependent dendritic integration and whether dendritic synaptic integration plays a role in PC computation.
@@
In this study, we employ detailed PC models that capture the intricacies of dendritic morphology to investigate dendritic nonlinear synaptic integration and its implications in information processing. Our examination of subthreshold synaptic summation in PCs reveals a significant sublinear synaptic integration profile. Notably, the strong sublinearity of spiny dendrites necessitates that PC have globally scattered inputs to fire somatic spikes. Furthermore, we observe that weak sublinear responses in smooth dendrites promote the generation of bursts, potentially leading to the generation of complex spikes. From the perspective of dendritic nonlinearity, this explains why the synaptic distribution in Purkinje cell is location-dependent. Importantly, this sublinearity enables a single neuron to compute feature binding problems. Our results suggest that while nonlinear dendrites may not supplant neurons as the fundamental computational units, they undoubtedly expand the computational capacities of neuronal systems.
@@ II-A Neuronal model
In this study, we employed a detailed neuron model incorporating neuromorphic structures to explore the dendritic nonlinear synaptic integration capabilities of Purkinje cells and how dendritic nonlinearity affects neuronal response and computational capabilities. This model extends the traditional Hodgkin-Huxley framework by including detailed representations of the neuron’s morphology, such as dendrites and soma compartments, to better capture the spatial and temporal dynamics of electrical signaling. The Hodgkin-Huxley model describes the membrane potential $\rm V$ through a set of nonlinear differential equations. These equations account for the ionic currents across the neuronal membrane, and leakage currents. In the neuromorphic model, these equations are applied to multiple compartments representing different parts of the neuron. The membrane potential dynamics for each compartment $\rm i$ are given by:
$\begin{split}C_{m}\frac{dV_{i}}{dt}=-I_{\text{ion},i}-I_{L,i}+I_{\text{syn},i}+\sum_{j}\frac{g_{ij}}{C_{m}}(V_{j}-V_{i})\end{split}$ (1)
@@
where $\rm C_{m}$ is the membrane capacitance, $\rm I_{\text{ion},i}$ is the ion current in compartment $\rm i$, $\rm I_{L,i}$ is the leakage current in compartment $\rm i$, $\rm I_{\text{syn},i}$ is the synaptic current applied to compartment $\rm i$, $\rm g_{ij}$ is the conductance between compartments $\rm i$ and $\rm j$, $\rm V_{j}$ is the membrane potential of compartment $\rm j$.
@@
Ion channels are fundamental components of the neuronal membrane, responsible for generating and propagating electrical signals in neurons. In this study, we modeled the behavior of these ion channels using the Hodgkin-Huxley framework. This framework provides a detailed biophysical description of how these ion channels contribute to the membrane potential dynamics. The ion current $\rm I_{\text{ion}}$ is given by:
$\begin{split}I_{\text{ion}}=g_{\text{ion}}m^{n}h^{i}(V-E_{\text{ion}})\end{split}$ (2)
@@
where: $\rm g_{\text{ion}}$ is the maximum ion conductance, $\rm m$ is the activation gating variable, $\rm h$is the inactivation gating variable, $\rm n$ and $\rm i$ represent the number of activation and inactivation gates respectively, $\rm V$ is the membrane potential, $\rm E_{\text{ion}}$ is the ion reversal potential. The gating variables $\rm m$ and $\rm h$ follow first-order kinetics with voltage-dependent rate constants:
$\begin{split}\frac{dm}{dt}=\alpha_{m}(1-m)-\beta_{m}m\\ \frac{dh}{dt}=\alpha_{h}(1-h)-\beta_{h}h\end{split}$ (3)
@@
where $\rm\alpha_{m,h}$ and $\rm\beta_{m,h}$ are rate constants for the ion channel. $\rm I_{L}$ is the leakage ion channel represents the passive flow of ions through the membrane, contributing to the resting membrane potential. The leakage current $\rm I_{L}$ is given by:
$\begin{split}I_{L}=g_{L}(V-E_{L})\end{split}$ (4)
where:$\rm g_{L}$ is the leakage conductance, $\rm V$ is the membrane potential, $\rm E_{L}$ is the leakage reversal potential.
@@
We used a 3D reconstruction of a mouse Purkinje cell available on the public archive www.neuromorpho.org (NMO-00865). The model contains a somatic compartment and 739 dendrites. The PC was modeled with the same parameters as before [41]. Membrane resistance ($\rm R_{m}$) and axial resistance ($\rm R_{i}$) were uniformly set to $\rm 5000\Omega/c{m^{2}}$ and $\rm 250\Omega/cm$, respectively. Membrane capacitance ($\rm C_{m}$) was set to $\rm 0.8F/c{m^{2}}$ in soma and mian dendrites, and $\rm 1.5F/c{m^{2}}$ in spiny dendrites. 13 different types of voltage-gated ion channels were installed in the PC model. Three ion channels (fast $\rm Na^{+}$ channel, persistent $\rm Na^{+}$ channel, anomalous rectifier channel) were solely added to the soma, and two ion channels (high-threshold $\rm Ca^{2+}$-activated ${\rm K^{+}}$channel, low-threshold $\rm Ca^{2+}$-activated ${\rm K^{+}}$ channel) were solely added to the dendrites. Eight of which (P-type $\rm C{a^{2+}}$ channel, T-type $\rm Ca^{2+}$ channel, class-E $\rm Ca^{2+}$ channel, persistent $\rm{K^{+}}$ channel, A-type $\rm{K^{+}}$ channel, D-type $\rm{K^{+}}$ channel, delayed rectifier, decay of sub-membrane $\rm Ca^{2+}$) were inserted into the soma and dendrites.
@@ II-B Synapse model
Synapses are critical structures that mediate communication between neurons through the release of neurotransmitters. The synapse models are based on the biophysical properties of synaptic currents and their interaction with the postsynaptic membrane. In this study, to explore the dendritic nonlinear synaptic integration ability of PCs, we established two excitatory synapse models: AMPA and NMDA. Excitatory synapses typically involve the release of neurotransmitters such as glutamate, which bind to receptors on the postsynaptic neuron, leading to depolarization. The excitatory postsynaptic current (EPSC) can be represented by a mathematical as:
$\begin{split}I_{syn}=g_{syn}*(V-E_{syn})\\ g_{syn}=g_{max}*w*Y*s\end{split}$ (5)
@@
$\rm{I_{syn}}$ is the receptor current, $\rm{g_{syn}}$ is the receptor conductance, $\rm V$ is the membrane potential, $\rm{{E_{syn}}}$ is the receptor reversal potential, $\rm{g_{\max}}$ is the maximum synaptic conductance, $\rm w$ is the connection weight between the synaptic stimulus and the neuron. NMDA-mediated currents often require AMPA-mediated depolarization to remove extracellular $\rm Mg^{2+}$ blockade of NMDA-associated channels, the associated channels open only when magnesium ions are blocked and NMDA receptors are activated. Therefore the scaling factor $\rm Y$ was:
$\begin{split}\left\{\begin{array}[]{l}Y=1/(1+0.24*exp(-154*v/0.027))~{}for~{}NMDA\\ Y=1~{}for~{}AMPA\end{array}\right.\end{split}$ (6)
The synaptic gating variable $\rm s$ follows double exponential kinetics:
$\begin{split}s(t)=e^{-(t-t_{\text{spike}})/\tau_{\text{decay}}}-e^{-(t-t_{\text{spike}})/\tau_{\text{rise}}}\end{split}$ (7)
@@
where $\rm\tau_{\text{rise}}$ is the rise time constant, $\rm\tau_{\text{decay}}$ is the decay time constant, and $\rm t_{spike}$ represents the time of the presynaptic spike, and rise time constant $\rm\tau_{rise}$ was set 0.5 ms, 8 ms for AMPA and NMDA, and decay time constant $\rm\tau_{decay}$ was set 5 ms, 30 ms for AMPA and NMDA, respectively. When the excitatory synapse includes both an AMPA and an NMDA component, they are co-located and always co-activated. When calculating dendritic nonlinearity, synapses are distributed starting at the 0.1 position of the dendrite, then added distally at intervals of $\rm d=0.2$ from the starting point.
@@
Each synapse received an independent spike train generated as a stimulus input. A single stimulation consists of a sequence of spikes containing spike times and inter-spike intervals, so we can generate a successive spike train by the previous spike plus the regular or irregular time intervals. In this study, we utilized a Poisson-distributed pulse sequence model to investigate the response of neurons to stochastic external stimuli. The spike train was generated according to a Poisson process, characterized by an average firing rate $\rm\lambda$ (pulses per second). The intervals between pulses, $\rm\Delta t$ , were exponentially distributed with parameter $\rm\lambda$ , ensuring that the pulse arrivals are memoryless and randomly timed. The probability density function of the inter-pulse intervals is given by:
$\begin{split}P(\Delta t)=\lambda e^{-\lambda\Delta t}\end{split}$ (8)
@@
where $\rm\Delta t$ is the time between consecutive pulses. Firstly, we generate a random number $\rm u$ uniformly distributed in the interval $(0,1)$. Then compute the inter-pulse interval $\rm\Delta t$ using the inverse transform sampling method:
$\begin{split}\Delta t=-\frac{1}{\lambda}\ln(1-u)\end{split}$ (9)
@@
and the next pulse arrival time is $\rm t_{i+1}=t_{i}+\Delta t$ and this process was repeated to generate a sequence of pulse arrival times over the desired simulation period.

### v2

@@ Abstract
A single neuron receives an extensive array of synaptic inputs through its dendrites, raising the fundamental question of how these inputs undergo integration and summation, culminating in the initiation of spikes in the soma. Experimental and computational investigations have revealed various modes of integration operations that include linear, superlinear, and sublinear summation. Interestingly, different types of neurons exhibit diverse patterns of dendritic integration depending on the spatial distribution of dendrites. The functional implications of these specific integration modalities remain largely unexplored. In this study, we employ the Purkinje cell (PC) as a model system to investigate these complex questions. Our findings reveal that PCs generally exhibit sublinear summation across their expansive dendrites. Both spatial and temporal input dynamically modulates the degree of sublinearity. Strong sublinearity necessitates the synaptic distribution in PCs to be globally scattered sensitive, whereas weak sublinearity facilitates the generation of complex firing patterns in PCs. Using dendritic branches characterized by strong sublinearity as computational units, we demonstrate that a neuron can successfully address the feature binding problem. Taken together, these results offer a systematic perspective on the functional role of dendritic sublinearity, inspiring a broader understanding of dendritic integration in various neuronal types.
@@ I Introduction
Neurons, the foundational computational entities within the complex brain system, precisely process signal input through their dendritic spine. The conventional understanding of dendritic function posits that these structures receive synaptic input and subsequently convey it to the soma, thus generating action potentials or spikes. This proposition suggests that a single neuron has computational capabilities to execute information processing through an extensive array of dendritic branches. However, the underlying integrating principles governing the diverse array of dendrites and neuron types remain elusive. In recent decades, experimental studies have facilitated laborious explorations of the integration of synaptic inputs along the dendritic domains of individual neurons, shedding light on the physiological basis of dendrite function [1, 2, 3, 4]. Dendrites have been shown to function as autonomous information processing units, capable of performing local computations, thus conferring substantial computational capabilities on single neurons [5, 6, 7, 8, 9, 10]. Furthermore, these insights have contributed significantly to our understanding of how neurons engage in communication with each other and the resultant impact on behavior [11, 12, 13, 14].
@@
The principal function of dendrites lies in the sophisticated integration and processing of synaptic inputs [6, 15]. Investigations focusing on pyramidal cells reveal that dendritic integration is characterized by superlinear dynamics, attributed to the activation of voltage-dependent channels and NMDARs [12, 16, 17, 18]. In particular, this superlinear integration has been observed primarily in cortical neurons [11, 19, 20], while sublinear summation has been documented in cerebellar interneurons  [21, 22]. The manifestation of dendritic nonlinearity allows neurons to function as feature detectors [14, 13], exerting influence on various aspects of brain function, including whisker sensation [23], orientation selectivity [24], sensory perception [25, 26], sensory-motor integration [23], and memory encoding [27, 28]. Furthermore, nonlinear dendritic integration is postulated to augment the computational capacity of neurons [29, 30], enabling them to compute linearly non-separable functions [31, 32]. The incorporation of nonlinear dendritic mechanisms into artificial neural networks is proposed to bring neurons in comparative alignment with their biological counterparts [33, 34, 28, 5]. The benefit of installing dendritic nonlinearity into artificial neural networks not only reduces power consumption and enhances accuracy [35, 36] but also mitigates communication costs within neural networks [37].
@@
Most investigations of dendritic nonlinearity have mainly focused on cortical neurons. However, there is a substantial gap in our understanding of the dendritic nonlinearity inherent in the most complex neuronal entities, namely cerebellar Purkinje cells (PCs) [38]. The complex dendritic structures of PCs contribute to the amplification of the computational repertoire within single neurons [39]. PCs, characterized by receiving thousands of synaptic inputs in a location-dependent manner, exhibit a distinctive synaptic organization, with the distal region of the dendritic tree receiving excitatory input from parallel fibers originating from granule cells [40, 41]. On the contrary, the proximal portion of the dendritic tree is innervated by the input of climbing fibers [42]. The interplay between synaptic input and dendritic nonlinear properties produces highly precise input-output functions [34]. Therefore, it is imperative to understand whether passive cable properties contribute to location-dependent dendritic integration and whether dendritic synaptic integration plays a role in PC computation.
@@
In this study, we employ detailed PC models that capture the intricacies of dendritic morphology to investigate dendritic nonlinear synaptic integration and its implications in information processing. Our examination of subthreshold synaptic summation in PCs reveals a significant sublinear synaptic integration profile. Notably, the strong sublinearity of spiny dendrites requires PCs to have globally scattered input to fire somatic spikes. Furthermore, we observe that weak sublinear responses in smooth dendrites promote the generation of bursts, potentially leading to the generation of complex spikes. From the perspective of dendritic nonlinearity, this explains why the synaptic distribution in PCs is location-dependent. Importantly, this sublinearity enables a single neuron to compute feature binding problems. Our results suggest that while nonlinear dendrites may not supplant neurons as the fundamental computational units, they undoubtedly expand the computational capacities of neuronal systems.
@@ II-A Neuronal model
In this study, we used a detailed neuron model incorporating neuromorphic structures to explore the dendritic nonlinear synaptic integration capabilities of Purkinje cells and how dendritic nonlinearity affects the neuronal response and computational capabilities. This model extends the traditional Hodgkin-Huxley framework [43] by including detailed representations of the neuron’s morphology, such as dendrites and soma compartments, to better capture the spatial and temporal dynamics of electrical signaling. The model describes the membrane potential $V$ through a set of nonlinear differential equations. These equations account for the ionic currents across the neuronal membrane and leakage currents. In the neuromorphic model, these equations are applied to multiple compartments representing different parts of the neuron. The membrane potential dynamics for each compartment $i$ are given by:
$\begin{split}C_{m,i}\frac{dV_{i}}{dt}=-I_{\text{ion},i}-I_{L,i}-I_{\text{syn},i}+\sum_{j}g_{ij}(V_{j}-V_{i}),\end{split}$ (1)
@@
where $C_{m,i}$ is the membrane capacitance of compartment $i$, $I_{\text{ion},i}$ is the ion current in compartment $i$, $I_{L,i}$ is the leakage current in compartment $i$, $I_{\text{syn},i}$ is the synaptic current applied to compartment $i$, $V_{j}$ is the membrane potential of compartment $j$, $g_{ij}$ is the conductance between compartments $i$ and $j$:
$\begin{split}g_{ij}=\frac{\pi d_{i}^{2}d_{j}^{2}}{2R_{i}\left(L_{i}d_{j}^{2}+L_{j}d_{i}^{2}\right)},\end{split}$ (2)
where $d_{i,j}$ and $L_{i,j}$ are the diameters and lengths of compartments $i$ and $j$, respectively. $R_{i}$ is the axial resistance.
@@
Ion channels are fundamental components of the neuronal membrane and are responsible for generating and propagating electrical signals in neurons. The ion current $I_{\text{ion}}$ is given by:
$\begin{split}I_{\text{ion}}=g_{\text{ion}}m^{n}h^{i}(V-E_{\text{ion}}),\end{split}$ (3)
@@
where $g_{\text{ion}}$ is the maximum ion conductance, $m$ is the activation gating variable, $h$ is the inactivation gating variable, $n$ and $i$ represent the number of activation and inactivation gates respectively, $V$ is the membrane potential, and $E_{\text{ion}}$ is the ion reversal potential. The gating variables $m$ and $h$ follow first-order kinetics with voltage-dependent rate constants:
$\begin{split}\frac{dm}{dt}=\alpha_{m}(1-m)-\beta_{m}m,\\ \frac{dh}{dt}=\alpha_{h}(1-h)-\beta_{h}h,\end{split}$ (4)
@@
where $\alpha_{m,h}$ and $\beta_{m,h}$ are rate constants for the ion channel. The leakage ion channel represents the passive flow of ions through the membrane, contributing to the resting membrane potential. The leakage current $I_{L}$ is given by:
$\begin{split}I_{L}&=g_{L}(V-E_{L}),\\ g_{L}&=1/R_{m},\end{split}$ (5)
where $g_{L}$ is the leakage conductance, $V$ is the membrane potential, $E_{L}$ is the leakage reversal potential, $R_{m}$ is the membrane resistance.
@@
We used a 3D reconstruction of a mouse PC (NMO-00865) available on the public archive (www.neuromorpho.org). The model contains a somatic compartment and 739 dendrites. The passive properties and voltage-dependent ionic channels used in this study are consistent with those reported in [44], which investigates the role of NMDA receptors across different dendritic branches and regions of Purkinje cells. PC model parameters were primarily derived from [45], with additional ionic channel parameters were adopted from [46, 47]. $R_{m}$ and $R_{i}$ were set to 5000 $\Omega*c{m^{2}}$ and 250 $\Omega*cm$, respectively. Membrane capacitance ($C_{m}$) was set to 0.8 $\mu F/c{m^{2}}$ in the soma and main dendrites, and 0.5 $\mu F/c{m^{2}}$ in spiny dendrites. 13 different types of voltage-gated ion channels were installed in the PC model. Three channels (fast $Na^{+}$ channel (NaF), persistent $Na^{+}$ channel (NaP), anomalous rectifier channel (Kh)) were solely added to the soma, and two channels (high-threshold $Ca^{2+}$-activated ${K^{+}}$channel (KC3), low-threshold $Ca^{2+}$-activated ${K^{+}}$ channel (K23)) were solely added to the dendrites. Eight (P-type $C{a^{2+}}$ channel (CaP), T-type $Ca^{2+}$ channel (CaT), class-E $Ca^{2+}$ channel (CaE), persistent ${K^{+}}$ channel (KM), A-type ${K^{+}}$ channel (KA), D-type ${K^{+}}$ channel (KD), delayed rectifier (Khh), decay of sub-membrane $Ca^{2+}$) were inserted into the soma and dendrites. The decay of sub-membrane $Ca^{2+}$ can be described as:
$\begin{split}&\frac{d[\text{Ca}^{2+}]_{i}}{dt}=\text{drive\_channel}+\text{drive\_pump}+\frac{[\text{Ca}^{2+}]_{\infty}-[\text{Ca}^{2+}]_{i}}{\tau_{r}},\\ &\text{drive\_channel}=-\frac{10000\cdot I_{Ca}}{2\cdot F\cdot\text{D}},\\ &\text{drive\_{pump}}=-\frac{k_{t}\cdot[\text{Ca}^{2+}]_{i}}{[\text{Ca}^{2+}]_{i}+k_{d}},\end{split}$ (6)
where the steady-state calcium concentration $[\text{Ca}^{2+}]_{\infty}$ is $4e^{-5}$ mM, the calcium removal time constant $\tau_{r}$ is 2ms, the Faraday constant (F) is 96,489, the depth of the shell (D) is $0.1\mu m$, the maximum rate of the pump $K_{t}$ is $4e^{-5}$ mM/ms, and the dissociation constant of the pump $k_{d}$ is $4e^{-5}$ mM. The parameters for the remaining ion channels are listed in Table  I.
TABLE I: Ion channel parameters.
Ion Channel Maximum Conductance (g_ion, nS) Activation Gate (n) Inactivation Gate (i) Reversal Potential (E_ion, mV)
NaF 10 3 1 45
NaP 0.001 3 \ 45
Kh 0.0005 1 \ -30
KC3 0.06 1 2 -85
K23 0.00039 1 2 -85
CaP 0.004 1 \ 135
CaI 0.0015 1 1 135
CaE 0.008 1 1 135
KM 0.00001 1 \ -85
KA 0.08 4 1 -85
KD 0.09 1 1 -85
Khh 0.0006 4 \ -85
@@ II-B Synapse model
Synapses are critical structures that mediate communication between neurons through the release of neurotransmitters. Synapse models are based on the biophysical properties of synaptic currents and their interaction with the postsynaptic membrane. To explore the dendritic nonlinear synaptic integration ability of PCs, two excitatory synapse models, AMPA and NMDA, were included. Excitatory synapses typically involve the release of neurotransmitters such as glutamate, which bind to receptors on the postsynaptic neuron, leading to depolarization. The excitatory postsynaptic current (EPSC) can be represented as:
$\begin{split}I_{syn}=g_{syn}*(V-E_{syn}),\\ g_{syn}=g_{max}*w*Y*S(t),\end{split}$ (7)
where ${I_{syn}}$ is the receptor current, ${g_{syn}}$ is the receptor conductance, $V$ is the membrane potential, ${{E_{syn}}}$ is the receptor reversal potential, ${g_{\max}}$ is the maximum synaptic conductance, and $w$ is the connection weight.
@@
NMDA-mediated currents typically require AMPA receptor-mediated depolarization to relieve the extracellular $Mg^{2+}$ blockade on NMDA-associated channels. These channels only open when the magnesium ions are displaced, allowing NMDA receptors to be activated concurrently. The scaling factor $Y$ is 1 for AMPA, while it is a nonlinear voltage ($V$) dependent function for NMDA: $Y=1/(1+0.24*exp(-154*V/0.027)).$
The synaptic gating variable $S(t)$ follows double exponential kinetics:
$\begin{split}S(t)=e^{-(t-t_{spike})/\tau_{decay}}-e^{-(t-t_{spike})/\tau_{rise}},\end{split}$ (8)
@@
where $\tau_{rise}$ is the rise time constant, $\tau_{decay}$ is the decay time constant, and $t_{spike}$ represents the time of the presynaptic spike. $\tau_{rise}$ was set to 0.5 ms for AMPA and 8 ms for NMDA. $\tau_{decay}$ was set to 5 ms for AMPA and 30 ms for NMDA [45, 48]. When the excitatory synapse includes both an AMPA and an NMDA component, they are co-located and always co-activated. When calculating dendritic nonlinearity, synapses are distributed starting at the 0.1 position of the dendrite, then added distally at intervals of $d=0.2$ from the starting point.
@@
Each synapse receives an independent spike train as stimulus input. A single stimulation consists of a sequence of spikes, including their spike times and inter-spike intervals. Consequently, a successive spike train can be generated by adding regular or irregular time intervals to the previous spike. We utilized a Poisson-distributed pulse sequence model to investigate the response of neurons to stochastic external stimuli. The spike train was generated according to a Poisson process, characterized by an average firing rate $\lambda$ (pulses per second). The intervals between pulses $\Delta t$ were exponentially distributed with parameter $\lambda$, ensuring that the pulse arrivals are memoryless and randomly timed. The probability density function of the inter-pulse intervals is given by:
$\begin{split}P(\Delta t)=\lambda e^{-\lambda\Delta t},\end{split}$ (9)
@@
where $\Delta t$ is the time between consecutive pulses. Firstly, we generate a random number $u$ uniformly distributed in the interval $(0,1)$. Then compute the inter-pulse interval $\Delta t$ using the inverse transform sampling method:
$\begin{split}\Delta t=-\frac{1}{\lambda}\ln(1-u),\end{split}$ (10)
@@
and the next pulse arrival time is $t_{i+1}=t_{i}+\Delta t$, and this process was repeated to generate a sequence of pulse arrival times over the desired simulation period.